OpenStudy (anonymous):

If someone would pretty please help me work this problem? :) solve. x^2=6x-6

OpenStudy (anonymous):

My answer choices are: x=3+_sqrt3 x=6+-sqrt3<=== I think this is it. I'm not completely sure though. x=3+-sqrt30 x=3+-sqrt6 I'm sorry, I don't know how to do the symbols :(

OpenStudy (anonymous):

\[x=3\pm \sqrt{3}\]

OpenStudy (anonymous):

I don't think this is it ^

OpenStudy (phi):

the first thing to do with this x^2=6x-6 is put it in standard form.

OpenStudy (phi):

add -6x +6 to both sides x^2 -6x +6=6x-6 -6x +6 or x^2 -6x +6=0

OpenStudy (anonymous):

This is it? \[x=6\pm \sqrt{3}\]

OpenStudy (phi):

take a peak at the video

OpenStudy (anonymous):

Okay :)